# G2RPO: Granular GRPO for Precise Reward in Flow Models

### v2

@@ Abstract
The integration of online reinforcement learning (RL) into diffusion and flow models has recently emerged as a promising approach for aligning generative models with human preferences. Stochastic sampling via Stochastic Differential Equations (SDE) is employed during the denoising process to generate diverse denoising directions for RL exploration. While existing methods effectively explore potential high-value samples, they suffer from sub-optimal preference alignment due to sparse and narrow reward signals. To address these challenges, we propose a novel Granular-GRPO (G2RPO) framework that achieves precise and comprehensive reward assessments of sampling directions in reinforcement learning of flow models. Specifically, a Singular Stochastic Sampling strategy is introduced to support step-wise stochastic exploration while enforcing a high correlation between the reward and the injected noise, thereby facilitating a faithful reward for each SDE perturbation. Concurrently, to eliminate the bias inherent in fixed-granularity denoising, we introduce a Multi-Granularity Advantage Integration module that aggregates advantages computed at multiple diffusion scales, producing a more comprehensive and robust evaluation of the sampling directions. Experiments conducted on various reward models, including both in-domain and out-of-domain evaluations, demonstrate that our G2RPO significantly outperforms existing flow-based GRPO baselines, highlighting its effectiveness and robustness.
11footnotetext: Equal contribution Corresponding author
@@ 1 Introduction
Recent advances in generative models, particularly diffusion models (Ho et al., 2020; Song et al., 2020a; b) and flow models (Lipman et al., 2022; Liu et al., 2022; Peebles & Xie, 2023), have revolutionized visual content creation, offering unprecedented capabilities in generating high-quality images (Rombach et al., 2022; Podell et al., 2023; Esser et al., 2024; Labs, 2024) and videos (Blattmann et al., 2023; Chen et al., 2024; Guo et al., 2023; Kong et al., 2024; Wan et al., 2025). However, a key challenge remains in aligning model outputs with the diverse and complex human preferences. To tackle this challenge, reinforcement learning from human feedback (RLHF) (Fan et al., 2023; Black et al., 2023) has emerged as a promising solution, characterized by its adaptability and cost-effectiveness. Paradigms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), Direct Policy Optimization (DPO) (Rafailov et al., 2023), and Group Relative Policy Optimization (GRPO) (Shao et al., 2024) have been introduced. Among these, GRPO stands out as an innovative online reinforcement learning approach. By leveraging group comparisons to optimize policies, GRPO eliminates the need for a separate value model, achieving greater flexibility and scalability.
Figure 1: Comparison between our G2RPO and existing studies. (a) G2RPO significantly outperforms DanceGRPO in reward scores (HPS-v2.1 in this figure). (b) Sampling strategy comparison. G2RPO acquire a dense reward by confining stochasticity to individual sampling steps. (c) Sampling grain comparison. G2RPO achieves a comprehensive evaluation of each sampling direction by integrating advantages from multi-granularity ODE denoising. (d) Visual Comparison. Compared to the baseline method, the images generated by G2RPO are more aligned with human preferences.
@@
To integrate GRPO into flow-based generative models, Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025) substitute the deterministic ODE sampler with an SDE formulation, wherein the injected stochasticity deliberately perturbs the denoising direction at each step. Although the resulting samples enable exhaustive per-step exploration for reinforcement learning, they simultaneously underscore the difficulty of attributing the final reward to any specific random perturbation, thereby constraining model trainability. Specifically, most existing flow-based GRPO methods encounter two core issues in evaluating group denoising directions: 1) Sparse reward: As shown in Fig. 1 (b), the final reward signal is uniformly assigned to each SDE sampling step, which cannot be precisely aligned with the sampling direction at each step, leading to inaccuracies for the optimization at individual steps. 2) Incomplete evaluation: As depicted in Fig. 1 (c), each denoising direction is bound to a fixed number of denoising steps, resulting in a singular granularity of denoised images, which impairs the reward model’s ability to conduct a comprehensive comparison across the group.
@@
To address these limitations, we propose Granular-GRPO (G2RPO), a novel online reinforcement learning framework specifically designed for precise and comprehensive reward signals. First, mirroring the sparse-reward problem (Hare, 2019; Liang et al., 2024) that plagues RLHF, the reward signal in the SDE sampling process is delivered only after an entire sequence of decisions. This long delay undermines the credit-assignment chain, preventing the linking of the terminal reward with any specific earlier action and thereby inducing sluggish, unstable learning. Therefore, we propose a simple yet effective sampling strategy, termed Singular Stochastic Sampling. As illustrated in Fig 1 (b), this strategy applies the SDE formulation at a single time step to generate a group of denoising directions, while employing deterministic ODE sampling for all other steps. By concentrating stochasticity at one specific step, the proposed method establishes a strong correlation between the reward signal and the injected noise, enabling stable model optimization. Secondly, we propose a Multi-Granularity Advantage Integration (MGAI) module. As depicted in Fig. 1(c), instead of binding each denoising direction to a fixed subsequent denoising granularity, the denoising directions in the same group are assigned to a spectrum of denoising steps, producing images with different granularities. The corresponding reward signals of these images are then fused into a unified advantage estimate, yielding a comprehensive evaluation of the current state’s value.
With the support of the Singular Stochastic Sampling strategy and the Multi-Granularity Advantage Integration module, G2RPO can provide a more precise and comprehensive reward signal, thereby enhancing the upper limit of the GRPO model training. As shown in Fig 1 (a), our reward curves exhibit stable and significant improvements over the baseline during training. Additionally, Fig 1 (d) illustrates the images generated by G2RPO, highlighting its advantages in text prompt adherence and detail fidelity.
Our contributions can be summarized as follows: (1) Granular-GRPO: A novel flow-based GRPO framework designed to provide a precise and comprehensive evaluation of the denoising directions sampled by the SDE, thereby improving the precision of model optimization. (2) Singular Stochastic Sampling: A sampling strategy confines stochasticity to individual sampling steps, addressing the sparse reward issue associated with long-range stochasticity injection. (3) Multi-Granularity Advantage Integration: A module integrates the advantages of multi-granularity denoised images and enables a comprehensive evaluation of each sampling direction. (4) Superior Performance: Extensive experiments across various reward models demonstrate that our G2RPO significantly outperforms existing baselines, demonstrating its effectiveness and robustness.
@@ 2 Related Work
Alignment for Large Language Models. Recent years have witnessed a paradigm shift from supervised fine-tuning (Dong et al., 2023; Sun, 2024) to multi-turn online reinforcement learning (Shani et al., 2024; Abdulhai et al., 2023) when aligning Large Language Models (LLMs) with human intent, which is known as Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2025; Ouyang et al., 2022). Early RLHF pipelines typically involve training a reward model from pairwise comparisons to predict human preferences and guide a policy model through reinforcement learning algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017). Despite their effectiveness, PPO introduces intensive computational overhead and is sensitive to reward model inaccuracies, motivating value-free alternatives such as Group Relative Policy Optimization (GRPO) (Shao et al., 2024). Adopted by leading LLMs including OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025), GRPO aims to optimize policies based on relative preferences within a group of samples, providing a robust signal for policy improvement, particularly when absolute rewards are difficult to define or noisy. These advancements in LLM alignment provide a strong foundation for exploring similar human-centric optimization strategies in visual generation domains.
@@
Alignment for Flow Models. Diffusion and Flow models (Ho et al., 2020; Song et al., 2020a; b; Peebles & Xie, 2023; Rombach et al., 2022), which offer flexible visual creation through an iterative denoising process, have revolutionized the field of visual synthesis and become a pivotal part of generative models. Building on the success of aligning LLMs with human preferences, similar techniques have recently been transplanted to diffusion and flow models (Podell et al., 2023; Esser et al., 2024; Labs, 2024). Pioneer works like DDPO (Black et al., 2023) and ReFL (Xu et al., 2023) apply PPO to finetune diffusion models for improved aesthetic performance and human feedback alignment. These methods face challenges inherent to RL, including high variance, low efficiency, and sparse reward. Diffusion-DPO (Wallace et al., 2024) adapts the Direct Preference Optimization (DPO) (Rafailov et al., 2023) framework to directly optimize diffusion models from paired preference data, bypassing the need for an explicit reward model but suffering from distribution shift since no new samples are collected during training. Recent efforts such as DanceGRPO (Xue et al., 2025) and Flow-GRPO (Liu et al., 2025) enable GRPO-style policy updates by converting the ODE sampling into an equivalent SDE to each timestep, thereby acquiring a group of denoising directions for statistical sampling and RL exploration for flow models. More recently, MixGRPO (Li et al., 2025) has improved training efficiency through a hybrid ODE-SDE sampling approach while maintaining comparable performance. However, these methods are generally constrained by sparse rewards due to long-range stochasticity injection and the binding of each sampling direction to a fixed denoising granularity. These paradigms restrict the ability to conduct a comprehensive evaluation of each sampling direction, limiting the optimization ceiling of GRPO training.
3 Preliminary
@@
For the flow-based GRPO methods (Xue et al., 2025; Liu et al., 2025; Li et al., 2025), the denoising process is first modeled as a multi-step Markov decision process (MDP). Given a prompt $\bm{c}$, the agent with a flow model $p_{\theta}$ produce a reverse-time trajectories defined as $\Gamma=\left(\mathbf{s}_{T},\mathbf{a}_{T},\mathbf{s}_{T-1},\mathbf{a}_{T-1},\ldots,\mathbf{s}_{0},\mathbf{a}_{0}\right)$, where $\mathbf{s}_{t}=(\bm{c},t,\bm{x}_{t})$ is the state at timestep $t$ and $\bm{x}_{t}$ is the corresponding noisy sample. Specifically, $\bm{x}_{T}\sim N(0,I)$ and $\bm{x}_{0}$ is the denoised image. The action $\mathbf{a}_{t}$ represents the single step denoising process with the policy $\pi_{\theta}$, indicating a sampling direction $\frac{d\bm{x}}{dt}$ from $\bm{x}_{t}$ to $\bm{x}_{t-1}$, i.e. $\bm{x}_{t-1}\sim\pi_{\theta}(\bm{x}_{t-1}|\bm{x}_{t},\bm{c})$.
@@
$d\bm{x}_{t}=\bm{v}_{\theta}(\bm{x}_{t},t)dt,$ (1)
where, $\bm{v}_{\theta}(\bm{x}_{t},t)$ is the model output, given the noisy sample $\bm{x}_{t}$ and timestep $t$.
To match GRPO’s stochastic sampling requirements, Flow-GRPO (Liu et al., 2025) converts the ODE into an equivalent SDE sampling with the same marginal distribution:
$d\bm{x}_{t}=\left(\bm{v}_{\theta}\left(\bm{x}_{t},t\right)+\frac{\sigma_{t}^{2}}{2t}\left(\bm{x}_{t}+(1-t)\bm{v}_{\theta}\left(\bm{x}_{t},t\right)\right)\right)dt+\sigma_{t}d\bm{w}_{t},$ (2)
where $d\bm{w}_{t}$ denotes Wiener process increments, and $\sigma_{t}$ controls the stochasticity injected into the sampling direction. Furthermore, it can be discretized via the Euler–Maruyama scheme:
$\bm{x}_{t+\Delta t}=\bm{x}_{t}+\left(\bm{v}_{\theta}\left(\bm{x}_{t},t\right)+\frac{\sigma_{t}^{2}}{2t}\left(\bm{x}_{t}+(1-t)\bm{v}_{\theta}\left(\bm{x}_{t},t\right)\right)\right)\Delta t+\sigma_{t}\sqrt{\Delta t}\epsilon,\quad\epsilon\sim\mathcal{N}(0,I).$ (3)
@@
GRPO Training. With SDE sampling, flow-based GRPO methods introduce stochasticity at each timestep to generate a group of $G$ images $\{\bm{x}_{0}^{i}\}^{G}_{i=1}$. Then the reward model assigns a score $R(\bm{x}^{i}_{0},\bm{c})$ to $\bm{x}_{0}^{i}$, and the advantage is computed as:
$A_{t}^{i}=\frac{R(\bm{x}_{0}^{i},\bm{c})-\operatorname{mean}(\{R(\bm{x}_{0}^{j},\bm{c})\}_{j=1}^{G})}{\operatorname{std}(\{R(\bm{x}_{0}^{j},\bm{c})\}_{j=1}^{G})}.$ (4)
@@
$\mathcal{J}_{\text{Flow-GRPO }}(\theta)=\mathbb{E}_{\bm{c}\sim\mathcal{C},\left\{\bm{x}^{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old }}}(\cdot\mid\bm{c})}f(r,A,\theta,\varepsilon,\beta),$ (5)
@@
$f(r,A,\theta,\varepsilon,\beta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\left(\min\left(r_{t}^{i}(\theta)A_{t}^{i},\operatorname{clip}\left(r_{t}^{i}(\theta),1-\varepsilon,1+\varepsilon\right)A_{t}^{i}\right)-\beta D_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right),$ (6)
$r_{t}^{i}(\theta)=\frac{p_{\theta}\left(\bm{x}_{t-1}^{i}\mid\bm{x}_{t}^{i},\bm{c}\right)}{p_{\theta_{\text{old }}}\left(\bm{x}_{t-1}^{i}\mid\bm{x}_{t}^{i},\bm{c}\right)}.$ (7)
Notably, $\beta$ is the hyperparameter that controls the proportion of KL loss. Following the practices of DanceGRPO and MixGRPO, we set $\beta=0$ to achieve a more stable training process.
4 Granular-GRPO
As an online RL algorithm, flow-based GRPO methods utilize SDE to sample a group of denoising directions for optimization. A core issue underlying this paradigm is to obtain precise and comprehensive assessments of each sampling direction. To this end, we introduce the G2RPO framework to (i) confine stochasticity to individual steps (Singular Stochastic Sampling) for more precise reward signals, and (ii) integrate the advantages derived from multi-granularity denoising results (Multi-Granularity Advantage Integration) to acquire a more comprehensive evaluation, as shown in Fig. 2.

### v3

# Fine-Grained GRPO for Precise Preference Alignment in Flow Models

@@ Abstract
The incorporation of online reinforcement learning (RL) into diffusion and flow-based generative models has recently gained attention as a powerful paradigm for aligning model behavior with human preferences. By leveraging stochastic sampling via Stochastic Differential Equations (SDEs) during the denoising phase, these models can explore a variety of denoising trajectories, enhancing the exploratory capacity of RL. However, despite their ability to discover potentially high-reward samples, current approaches often struggle to effectively align with preferences due to the sparsity and narrowness of reward feedback. To overcome this limitation, we introduce a novel framework called Granular-GRPO (G2RPO), which enables fine-grained and comprehensive evaluation of sampling directions in the RL training of flow models. Specifically, we propose a Singular Stochastic Sampling mechanism that supports step-wise stochastic exploration while ensuring strong correlation between injected noise and reward signals, enabling more accurate credit assignment to each SDE perturbation. Additionally, to mitigate the bias introduced by fixed-granularity denoising, we design a Multi-Granularity Advantage Integration module that aggregates advantages computed across multiple diffusion scales, resulting in a more robust and holistic assessment of sampling trajectories. Extensive experiments on various reward models, including both in-domain and out-of-domain settings, demonstrate that our G2RPO outperforms existing flow-based GRPO baselines, highlighting its effectiveness and generalization capability.
Figure 1: Comparison between our G2RPO and existing studies. (a) G2RPO outperforms DanceGRPO in reward scores (HPS-v2.1 in this figure). (b) Sampling strategy comparison. G2RPO acquire a dense reward by confining stochasticity to individual sampling steps. (c) Sampling grain comparison. G2RPO achieves a comprehensive evaluation of each sampling direction by integrating advantages from multi-granularity ODE denoising. (d) Visual Comparison. Compared to the baseline method, the images generated by G2RPO are more aligned with human preferences.
11footnotetext: Equal contribution. Corresponding authors.
@@ 1 Introduction
Recent advances in generative models, particularly diffusion models [ho2020denoising, song2020denoising, song2020score] and flow models [lipman2022flow, liu2022flow, peebles2023scalable], have revolutionized visual content creation, offering unprecedented capabilities in generating high-quality images [rombach2022high, podell2023sdxl, esser2024scaling, flux2024] and videos [blattmann2023stable, chen2024videocrafter2, guo2023animatediff, kong2024hunyuanvideo, wan2025wan]. However, a key challenge remains in aligning model outputs with the diverse and complex human preferences. To tackle this challenge, reinforcement learning from human feedback (RLHF) [fan2023dpok, black2023training] has emerged as a promising solution, characterized by its adaptability and cost-effectiveness. Paradigms such as Proximal Policy Optimization (PPO) [schulman2017proximal], Direct Policy Optimization (DPO) [rafailov2023direct], and Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] have been introduced. Among these, GRPO stands out as an innovative online reinforcement learning approach. Leveraging group comparisons to optimize policies, GRPO eliminates the need for a separate value model, achieving greater flexibility and scalability.
@@
To integrate GRPO into flow-based generative models, Flow-GRPO [liu2025flow] and DanceGRPO [xue2025dancegrpo] substitute the deterministic ODE sampler with an SDE formulation, wherein the injected stochasticity deliberately perturbs the denoising direction at each step. Although the resulting samples enable exhaustive per-step exploration for reinforcement learning, they simultaneously underscore the difficulty of attributing the final reward to any specific random perturbation, thereby constraining model trainability. Specifically, most existing flow-based GRPO methods encounter two core issues in evaluating group denoising directions: 1) Sparse reward: As shown in Fig. 1 (b), the final reward signal is uniformly assigned to each SDE sampling step, which cannot be precisely aligned with the sampling direction at each step, leading to inaccuracies for the optimization at individual steps. 2) Incomplete evaluation: While the same single-step SDE sampling direction can yield different rewards under varying denoising intervals, an ideal high-quality direction (Sample 2 and Sample 3 shown in Fig. 3) intuitively is robust to the granularity of the remaining sampling trajectory. This means that although the absolute quality of the final results may vary with the number of steps, an excellent denoising direction still maintains a superior relative position within the group. Nevertheless, as shown in Fig. 1(c), existing methods suffer from the drawback that each denoising direction is rigidly associated with a fixed number of denoising steps. Consequently, it leads to a single granularity of the denoised images, making it impossible to conduct a thorough comparison across the group.
@@
To address these limitations, we propose Granular-GRPO (G2RPO), a novel online reinforcement learning framework specifically designed for precise and comprehensive reward signals. First, mirroring the sparse-reward problem [hare2019dealing, liang2024step] that plagues RLHF, the reward signal in the SDE sampling process is delivered only after an entire sequence of decisions. This long delay undermines the credit-assignment chain, preventing the linking of the terminal reward with any specific earlier action and thereby inducing sluggish, unstable learning. Therefore, we propose a simple yet effective sampling strategy, termed Singular Stochastic Sampling. As illustrated in Fig 1 (b), this strategy applies the SDE formulation at a single time step to generate a group of denoising directions, while employing deterministic ODE sampling for all other steps. By concentrating stochasticity at one specific step, the proposed method establishes a strong correlation between the reward signal and the injected noise, enabling stable model optimization. Secondly, we propose a Multi-Granularity Advantage Integration (MGAI) module. As depicted in Fig. 1(c), instead of binding each denoising direction to a fixed subsequent denoising granularity, the denoising directions in the same group are assigned to a spectrum of denoising steps, producing images with different granularities. The corresponding reward signals of these images are then fused into a unified advantage estimate, yielding a comprehensive evaluation of the current state’s value. With the support of the Singular Stochastic Sampling strategy and the Multi-Granularity Advantage Integration module, G2RPO can provide a more precise and comprehensive reward signal, thereby enhancing the upper limit of the GRPO model training. As shown in Fig 1 (a), our reward curves exhibit stable improvements over the baseline during training. Additionally, Fig 1 (d) illustrates the images generated by G2RPO, highlighting its advantages in text prompt adherence and detail fidelity.
Our contributions can be summarized as follows:
  • Granular-GRPO: A novel flow-based GRPO framework designed to provide a precise and comprehensive evaluation of the denoising directions sampled by the SDE, thereby improving the precision of model optimization.
  • Singular Stochastic Sampling: A sampling strategy confines stochasticity to individual sampling steps, addressing the sparse reward issue associated with long-range stochasticity injection.
  • Multi-Granularity Advantage Integration: A module integrates the advantages of multi-granularity denoised images and enables a comprehensive evaluation of each sampling direction.
@@ 2.1 Alignment for Large Language Models
Recent years have witnessed a paradigm shift from supervised fine-tuning [dong2023abilities, sun2024supervised] to online reinforcement learning [shani2024multi, abdulhai2023lmrl, jaech2024openai] when aligning Large Language Models (LLMs) with human intent, which is known as Reinforcement Learning from Human Feedback (RLHF) [bai2025qwen2, ouyang2022training]. Early RLHF pipelines typically involve training a reward model from pairwise comparisons to predict human preferences and guide a policy model through reinforcement learning algorithms like Proximal Policy Optimization (PPO) [schulman2017proximal]. Despite their effectiveness, PPO introduces intensive computational overhead due to its iterative optimization process and the need for frequent interactions with the environment. To address these computational challenges, value-free alternatives have been proposed, such as Group Relative Policy Optimization (GRPO) [shao2024deepseekmath]. Adopted by leading LLM DeepSeek-R1 [guo2025deepseek], GRPO aims to optimize policies based on relative preferences within a group of samples, providing a robust signal for policy improvement, particularly when absolute rewards are difficult to define or noisy. These advancements in LLM alignment provide a strong foundation for exploring similar human-centric optimization strategies in visual generation domains.
2.2 Alignment for Flow Models
@@
Diffusion and Flow models [ho2020denoising, song2020denoising, song2020score, peebles2023scalable, rombach2022high], which offer flexible visual creation through an iterative denoising process, have revolutionized the field of visual synthesis and become a pivotal part of generative models. Building on the success of aligning LLMs with human preferences, similar techniques have recently been transplanted to diffusion and flow models [podell2023sdxl, esser2024scaling, flux2024]. Pioneer works like DDPO [black2023training] and ReFL [xu2023imagereward] apply PPO to finetune diffusion models for improved aesthetic performance and human feedback alignment. These methods face challenges inherent to RL, including high variance, low efficiency, and sparse reward. Diffusion-DPO [wallace2024diffusion] adapts the Direct Preference Optimization (DPO) [rafailov2023direct] framework to directly optimize diffusion models from paired preference data, bypassing the need for an explicit reward model but suffering from distribution shift since no new samples are collected during training. Recent efforts such as DanceGRPO [xue2025dancegrpo] and Flow-GRPO [liu2025flow] enable GRPO-style policy updates by converting the ODE sampling into an equivalent SDE to each timestep, thereby acquiring a group of denoising directions for statistical sampling and RL exploration for flow models. More recently, MixGRPO [li2025mixgrpo] has improved training efficiency through a hybrid ODE-SDE sampling approach while maintaining comparable performance. However, these methods are generally constrained by sparse rewards due to long-range stochasticity injection and the binding of each sampling direction to a fixed denoising granularity. These paradigms restrict the ability to conduct a comprehensive evaluation of each sampling direction, limiting the optimization ceiling of GRPO training.
3 Granular-GRPO
As an online RL algorithm, flow-based GRPO methods utilize ODE-to-SDE conversion with the original model’s marginal distribution to sample a group of denoising directions for optimization. A core issue underlying this paradigm is to obtain precise and comprehensive assessments of each sampling direction. To this end, we introduce the Granular-GRPO framework to (i) confine stochasticity to individual steps (Singular Stochastic Sampling) for more precise reward signals, and (ii) integrate the advantages derived from multi-granularity denoising results (Multi-Granularity Advantage Integration) to acquire a more comprehensive evaluation, as shown in Fig. 2.
3.1 Preliminary
@@
For the flow-based GRPO methods [xue2025dancegrpo, liu2025flow, li2025mixgrpo], the denoising process is first modeled as a multi-step Markov decision process (MDP). Given a prompt $\boldsymbol{c}$, the agent with a flow model $p_{\theta}$ produce a reverse-time trajectories defined as $\Gamma=\left(\mathbf{s}_{T},\mathbf{a}_{T},\mathbf{s}_{T-1},\mathbf{a}_{T-1},\ldots,\mathbf{s}_{0},\mathbf{a}_{0}\right)$, where $\mathbf{s}_{t}=(\boldsymbol{c},t,\boldsymbol{x}_{t})$ is the state at timestep $t$ and $\boldsymbol{x}_{t}$ is the corresponding noisy sample. Specifically, $\boldsymbol{x}_{T}\sim N(0,I)$ and $\boldsymbol{x}_{0}$ is the denoised image. The action $\mathbf{a}_{t}$ represents the single step denoising process with the policy $\pi_{\theta}$, indicating a sampling direction $\frac{d\boldsymbol{x}}{dt}$ from $\boldsymbol{x}_{t}$ to $\boldsymbol{x}_{t-1}$, i.e. $\boldsymbol{x}_{t-1}\sim\pi_{\theta}(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{c})$.
@@
$d\boldsymbol{x}_{t}=\boldsymbol{v}_{\theta}(\boldsymbol{x}_{t},t)dt,$ (1)
where, $\boldsymbol{v}_{\theta}(\boldsymbol{x}_{t},t)$ is the model output, given the noisy sample $\boldsymbol{x}_{t}$ and timestep $t$.
To match GRPO’s stochastic sampling requirements, Flow-GRPO [liu2025flow] converts the ODE into an equivalent SDE sampling with the same marginal distribution:
$d\boldsymbol{x}_{t}=(\boldsymbol{v}_{\theta}(\boldsymbol{x}_{t},t)+\frac{\sigma_{t}^{2}}{2t}(\boldsymbol{x}_{t}+(1-t)\boldsymbol{v}_{\theta}(\boldsymbol{x}_{t},t)))dt+\sigma_{t}d\boldsymbol{w}_{t},$ (2)
where $d\boldsymbol{w}_{t}$ denotes Wiener process increments, and $\sigma_{t}$ controls the stochasticity injected into the sampling direction. Furthermore, it can be discretized via the Euler–Maruyama scheme:
$\boldsymbol{x}_{t+\Delta t}$ $=\boldsymbol{x}_{t}+\bigl(\boldsymbol{v}_{\theta}(\boldsymbol{x}_{t},t)+\tfrac{\sigma_{t}^{2}}{2t}(\boldsymbol{x}_{t}+(1-t)\boldsymbol{v}_{\theta}(\boldsymbol{x}_{t},t))\bigr)\Delta t$
$\quad+\sigma_{t}\sqrt{\Delta t}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I).$ (3)
@@
GRPO Training. With SDE sampling, flow-based GRPO methods introduce stochasticity at each timestep to generate a group of $G$ images $\{\boldsymbol{x}_{0}^{i}\}^{G}_{i=1}$. Then the reward model assigns a score $R(\boldsymbol{x}^{i}_{0},\boldsymbol{c})$ to $\boldsymbol{x}_{0}^{i}$, and the advantage is computed as:
$A_{t}^{i}=\frac{R(\boldsymbol{x}_{0}^{i},\boldsymbol{c})-\operatorname{mean}(\{R(\boldsymbol{x}_{0}^{j},\boldsymbol{c})\}_{j=1}^{G})}{\operatorname{std}(\{R(\boldsymbol{x}_{0}^{j},\boldsymbol{c})\}_{j=1}^{G})}.$ (4)
@@
$\mathcal{J}_{\text{Flow-GRPO }}(\theta)=\mathbb{E}_{\boldsymbol{c}\sim\mathcal{C},\left\{\boldsymbol{x}^{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old }}}(\cdot\mid\boldsymbol{c})}f(r,A,\theta,\varepsilon,\beta),$ (5)
@@
$f(r,A,\theta,\varepsilon,\beta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}(\min(r_{t}^{i}(\theta)A_{t}^{i},$
$\operatorname{clip}(r_{t}^{i}(\theta),1-\varepsilon,1+\varepsilon)A_{t}^{i})-\beta D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}}),$ (6)
$r_{t}^{i}(\theta)=\frac{p_{\theta}\left(\boldsymbol{x}_{t-1}^{i}\mid\boldsymbol{x}_{t}^{i},\boldsymbol{c}\right)}{p_{\theta_{\text{old }}}\left(\boldsymbol{x}_{t-1}^{i}\mid\boldsymbol{x}_{t}^{i},\boldsymbol{c}\right)}.$ (7)
Notably, $\beta$ is the hyperparameter that controls the proportion of KL loss. Following the practices of DanceGRPO and MixGRPO, $\beta=0$ to achieve a stable training process.